# Stimulated magnon scattering by non-degenerate parametric excitation

### v1

@@ Abstract
Nonlinear spin wave dynamics in magnetic thin films and devices has found renewed interest due to potential applications in unconventional computing. [1, 2, 3, 4] For neuro-inspired computing, a promising approach involves mapping the reciprocal space ($k$-space) spanned by spin wave eigenmodes to a recurrent neural network, whereby information propagates through the network by nonlinear processes such as magnon scattering. For example, three-magnon scattering in vortex states within micron-sized disks has proven effective for pattern recognition tasks, underpinning the nonlinear transformation between input and output magnon spectra in the radiofrequency (RF) domain. [4]
@@
We study the magnon population dynamics using mode-filtering. [17] This approach involves analysing the magnetization fluctuations $\delta\mathbf{m}(\mathbf{r},t)$ about the ground state $\mathbf{m}_{0}(\mathbf{r})$, $\mathbf{m}(\mathbf{r},t)=\mathbf{m}_{0}(\mathbf{r})+\delta\mathbf{m}(\mathbf{r},t)$, whereby the fluctuations are expressed in terms of normal modes
$\delta\mathbf{m}(\mathbf{r},t)\simeq\sum_{\kappa=1}^{\infty}\mathbf{a}_{\kappa}(\mathbf{r},t)\psi_{\kappa}(\mathbf{r}).$ (3)
The scalar fields $\psi_{\kappa}(\mathbf{r})$ represent spatial mode profiles and $\mathbf{a}_{\kappa}(\mathbf{r},t)$ the mode amplitudes that are defined in the plane perpendicular to the static magnetization, $\mathbf{m}_{0}$. The corresponding mode population is given by $n_{\kappa}\simeq\|\mathbf{a}_{\kappa}\|^{2}$[17] By projecting out these amplitudes directly in the simulations, we can track the time evolution of individual modes even when the frequency spacing between them are below the resolution allowable by the finite duration of the simulations. With this method, we can unambiguously identify which modes take part in scattering processes, offering an advantage over frequency-domain techniques since phenomena like frequency pulling and mutual phase locking can complicate spectral analyses. [17] The normal modes are obtained from diagonalizing a dynamical matrix using the method discussed in Refs. 18, 19, 20 and implemented in the open-source code magnum.np[21]
@@
Fig. 1(d) illustrates the mode-resolved method by presenting the data in Fig. 1(c) in terms of the mode populations instead. $n_{\kappa}$ for the first 200 modes are recorded at the end of the simulation, $t=t_{f}=500$ ns, and plotted using a color map as a function of the input frequency. Like in Fig. 1(c), we can identify intervals in $\nu_{\mathrm{rf}}$ in which pairs of modes $(\kappa_{1},\kappa_{2})$ are excited by the driving field (typically in the range $\kappa<80$), in additional to modes that directly couple to the RF field in the range $\kappa>140$. We also observe a dark horizontal line corresponding to $\kappa=12$, the quasi-uniform mode at this value of $B_{0}$, which indicates that it is strongly driven non-resonantly by $b_{\mathrm{rf}}$.
@@
Figure 2: (a) Selected mode profiles. (b) PSD under a driving frequency of $\nu_{\mathrm{rf}}=6.90$ GHz. (c) PSD under a driving frequency of $\nu_{\mathrm{rf}}=7.32$ GHz. (d) Mode population dynamics corresponding to (b). (e) Mode population dynamics corresponding to (c).
The power spectral density (PSD) is shown in Fig. 2(b) for two frequency ranges of interest. The top panel of Fig. 2(b) gives the range encompassing $\nu_{\mathrm{rf}}$ where peaks of scattered modes can be seen around $\nu_{\mathrm{rf}}/2$. The gray curve represents the contribution from $\kappa=12$, which exhibits the strongest response at $\nu_{\mathrm{rf}}$, confirming non-resonant excitation by the RF field, along with a weaker response around 2.9 GHz, which corresponds to its natural frequency. The bottom panel of Fig. 2(b) gives a magnified view of the scattered mode spectra, where peaks are labeled by $\kappa$. $\kappa=28$ and 30 are excited simultaneously as their natural frequencies are within 1 MHz of each other. The transient dynamics of representative scattered modes, along with $\kappa=12$, is given in Fig. 2(d). $\kappa=12$ is driven to steady state within tens of ns, while one of the primary scattered modes, $\kappa=39$, exhibits an exponential growth over around 200 ns before reaching steady state. Secondary pairs of scattered modes, represented by $\kappa=44$ and 48, exhibit an incubation time of over 100 ns before growing exponentially and saturating, while remaining orders of magnitude weaker than the primary scattered mode but comparable to $\kappa=12$. This incubation delay is similar to behavior observed for parallel pumping in the same geometry. [17] Above the curves in Fig. 2(d), we use a color map to present the same time dependence of the $n_{\kappa}$ on a logarithmic scale. We will rely heavily on this visualization scheme in subsequent figures.
A similar analysis of the scattered mode PSD and population dynamics is presented in Figs. 2(c) and 2(e), respectively, for a driving frequency of $\nu_{\mathrm{rf}}=7.32$ GHz. In contrast to the previous case, only one pair of scattered modes is observed, $\kappa=39$ and 48. As Fig. 2(e) shows, their population dynamics and levels mirror one another, as expected for the scattering process in Fig. 1(a). Note that the populations shown in Figs. 2(d) and 2(e) are orders of magnitude above their thermal levels, which lie in the range of $n=10^{-3}$ to $10^{-2}$.
@@
Figure 3: Population dynamics under dual excitation with $\nu_{A}=6.90$ GHz and $\nu_{B}=7.32$ GHz. (a) $\nu_{A}\rightarrow\nu_{A}+\nu_{B}$. (b) $\nu_{B}\rightarrow\nu_{B}+\nu_{A}$. (c) Non-degenerate scattering resulting from $\nu_{A}$ and $\nu_{B}$ separately. (d) Stimulated magnon scattering resulting in $\kappa=59$, with its profile shown in the inset. Numbers in parenthesis in (c) and (d) refer to the mode index $\kappa$.
@@
Similar stimulated processes also occur for dual excitations that do not share any common scattering channels. An example is given in Fig. 4 for $\nu_{A}=6.90$ GHz and $\nu_{B}=7.51$ GHz.
Figure 4: Population dynamics under dual excitation with $\nu_{A}=6.90$ GHz and $\nu_{B}=7.51$ GHz. (a) $\nu_{A}\rightarrow\nu_{A}+\nu_{B}$. (b) $\nu_{B}\rightarrow\nu_{B}+\nu_{A}$. (c) Non-degenerate scattering resulting from $\nu_{A}$ and $\nu_{B}$ separately. (d) Stimulated magnon scattering resulting in $\kappa=59$, with its profile shown in the inset. Numbers in parenthesis in (c) and (d) refer to the mode index $\kappa$.
@@
Figure 5(c) displays the scattering matrix for the ellipse. Similar features to the disk are seen, such as the variability in the number of stimulated modes and the non-commutativity of the excitation pulses, but there are also some striking differences, such as the distribution of $(\nu_{A},\nu_{B})$ leading to stimulated generation. Specifically, these are concentrated along well-defined rows or columns in the matrix plot, suggesting that only a handful of scattered pairs, i.e., involving 5.78 and 8.50 GHz in particular, are associated with silent channels most amenable to cross-stimulation. Finally, Fig. 5(d) presents the scattering matrix for the rectangle. A striking difference here is the large number of combinations that result in stimulated generation. These results show that stimulated magnon scattering can appear in a wide variety of geometries, but its prevalence likely depends on details related to the spectrum of the normal modes and coupling efficiency to the driving fields.
Based on these results, we envisage the following avenues for future study. From the theoretical perspective, it would be interesting to explore whether the stimulated processes relate to three-magnon splitting of the off-resonantly driven uniform mode, or to four-magnon processes involving recombination of parametrically-excited modes. This could be quantified by examining coupled equations of motion for the mode amplitudes, which could be constructed from spin wave theory. [19, 20] From the experimental perspective, the prediction of qualitatively different scattering matrices for different geometrical shapes could be tested directly using micro-focus Brillouin light scattering, which has already been employed to reveal non-degenerate parametric excitation in thin-film discs. [10] On a broader level, such stimulated processes could also be investigated for nonuniform magnetic ground states such as vortices.
@@ Acknowledgements.
We thank M. Massouras and G. de Loubens for fruitful discussions. This work was supported by the Horizon 2020 Framework Programme of the European Commission under contract number 899646 (k-Net).
@@
The data supporting the findings of this study are available from the corresponding author upon reasonable request. We used scientific color maps developed by Crameri et al., [23] namely batlowW, oslo, and vik, in order to minimize visual distortion of the data and artifacts for readers with color-vision deficiencies. [24]

### v2

@@ Abstract
Nonlinear spin wave dynamics in magnetic thin films and devices has found renewed interest due to potential applications in unconventional computing. [1, 2, 3, 4] For neuro-inspired computing, a promising approach involves mapping the reciprocal space ($k$-space) spanned by spin wave eigenmodes to a recurrent neural network, whereby information propagates through the network by nonlinear processes such as magnon scattering. For example, three-magnon scattering in vortex states within micrometer-sized disks has proven effective for pattern recognition tasks, underpinning the nonlinear transformation between input and output magnon spectra in the radio frequency (RF) domain. [4]
@@
We study the magnon population dynamics using mode-projection. [17] This approach involves analyzing the magnetization fluctuations $\delta\mathbf{m}(\mathbf{r},t)$ about the ground state $\mathbf{m}_{0}(\mathbf{r})$, $\mathbf{m}(\mathbf{r},t)=\mathbf{m}_{0}(\mathbf{r})+\delta\mathbf{m}(\mathbf{r},t)$, whereby the fluctuations are expressed in terms of normal modes
$\delta\mathbf{m}(\mathbf{r},t)\simeq\sum_{\kappa=1}^{\infty}c_{\kappa}(t)\bm{\psi}_{\kappa}(\mathbf{r})+\mathrm{c.c.},$ (3)
where “c.c.” means complex conjugate. The vector fields $\bm{\psi}_{\kappa}(\mathbf{r})$ represent normal mode profiles and $c_{\kappa}(t)$ the complex mode amplitudes, such that the corresponding mode population is given by $n_{\kappa}=c^{*}_{\kappa}c_{\kappa}$. By projecting out these amplitudes directly in the simulations, we can track the time evolution of individual modes even when the frequency spacing between them is below the resolution allowable by the finite duration of the simulations. With this method, we can unambiguously identify which modes take part in scattering processes, offering an advantage over frequency-domain techniques since phenomena like frequency pulling and mutual phase locking can complicate spectral analyses. [17] The normal modes are obtained from diagonalizing a dynamical matrix using the method discussed in Refs. 18, 19, 20 and implemented in the open-source code magnum.np[21]
@@
Figure 1(d) illustrates the mode-resolved method by presenting the data in Fig. 1(c) in terms of the mode populations instead. $n_{\kappa}$ for the first 200 modes are recorded at the end of the simulation, $t=t_{f}=500$ ns, and plotted using a color map as a function of the input frequency. Like in Fig. 1(c), we can identify intervals in $\nu_{\mathrm{rf}}$ in which pairs of modes $(\kappa_{1},\kappa_{2})$ are excited by the driving field (typically in the range $\kappa<80$), in addition to modes that directly couple to the RF field in the range $\kappa>140$. We also observe a dark horizontal line corresponding to $\kappa=12$, the quasi-uniform mode at this value of $B_{0}$, which indicates that it is strongly driven non-resonantly by $b_{\mathrm{rf}}$.
@@
Figure 2: (a) Selected mode profiles, $\psi_{\kappa}\equiv\bm{\psi}_{\kappa}\cdot\hat{\mathbf{z}}$. (b) PSD under a driving frequency of $\nu_{\mathrm{rf}}=6.90$ GHz. (c) PSD under a driving frequency of $\nu_{\mathrm{rf}}=7.32$ GHz. (d) Mode population dynamics corresponding to (b). (e) Mode population dynamics corresponding to (c).
The power spectral density (PSD) is shown in Fig. 2(b) for two frequency ranges of interest. The top panel of Fig. 2(b) gives the range encompassing $\nu_{\mathrm{rf}}$ where peaks of scattered modes can be seen around $\nu_{\mathrm{rf}}/2$. The gray curve represents the contribution from $\kappa=12$, which exhibits the strongest response at $\nu_{\mathrm{rf}}$, confirming non-resonant excitation by the RF field, along with a weaker response around 2.9 GHz, which corresponds to its natural frequency. The bottom panel of Fig. 2(b) gives a magnified view of the scattered mode spectra, where peaks are labeled by $\kappa$. $\kappa=28$ and 30 are excited simultaneously as their natural frequencies are within 1 MHz of each other. The transient dynamics of representative scattered modes, along with $\kappa=12$, is given in Fig. 2(d). $\kappa=12$ is driven to steady state within tens of nanoseconds, while one of the primary scattered modes, $\kappa=39$, exhibits an exponential growth over around 200 ns before reaching steady state. Secondary pairs of scattered modes, represented by $\kappa=44$ and 48, exhibit an incubation time of over 100 ns before growing exponentially and saturating, while remaining orders of magnitude weaker than the primary scattered mode but comparable to $\kappa=12$. This incubation delay is similar to the behavior observed for parallel pumping in the same geometry. [17] Above the curves in Fig. 2(d), we use a color map to present the same time dependence of $n_{\kappa}$ on a logarithmic scale. We will rely heavily on this visualization scheme in subsequent figures.
A similar analysis of the scattered mode PSD and population dynamics is presented in Figs. 2(c) and 2(e), respectively, for a driving frequency of $\nu_{\mathrm{rf}}=7.32$ GHz. In contrast to the previous case, only one pair of scattered modes is observed, $\kappa=39$ and 48. As Fig. 2(e) shows, their population dynamics and levels mirror one another, as expected for the scattering process in Fig. 1(a). Note that the populations shown in Figs. 2(d) and 2(e) are orders of magnitude above their thermal levels, which lie in the range of $n=10^{-3}$- $10^{-2}$.
@@
Figure 3: Population dynamics under dual excitation with $\nu_{A}=6.90$ and $\nu_{B}=7.32$ GHz. (a) $\nu_{A}\rightarrow\nu_{A}+\nu_{B}$. (b) $\nu_{B}\rightarrow\nu_{B}+\nu_{A}$. (c) Non-degenerate scattering resulting from $\nu_{A}$ and $\nu_{B}$ separately. (d) Stimulated magnon scattering resulting in $\kappa=59$, with its profile shown in the inset. Numbers in parenthesis in (c) and (d) refer to the mode index $\kappa$.
@@
Similar stimulated processes also occur for dual excitations that do not share any common scattering channels. An example is given in Fig. 4 for $\nu_{A}=6.90$ and $\nu_{B}=7.51$ GHz.
Figure 4: Population dynamics under dual excitation with $\nu_{A}=6.90$ and $\nu_{B}=7.51$ GHz. (a) $\nu_{A}\rightarrow\nu_{A}+\nu_{B}$. (b) $\nu_{B}\rightarrow\nu_{B}+\nu_{A}$. (c) Non-degenerate scattering resulting from $\nu_{A}$ and $\nu_{B}$ separately. (d) Stimulated magnon scattering resulting in $\kappa=59$, with its profile shown in the inset. Numbers in parenthesis in (c) and (d) refer to the mode index $\kappa$.
@@
Figure 5(c) displays the scattering matrix for the ellipse. Similar features to the disk are seen, such as the variability in number of stimulated modes and the non-commutativity of the excitation pulses, but there are also some striking differences, such as the distribution of $(\nu_{A},\nu_{B})$ leading to stimulated generation. Specifically, these are concentrated along well-defined rows or columns in the matrix plot, suggesting that only a handful of scattered pairs, i.e., involving 5.78 and 8.50 GHz in particular, are associated with silent channels most amenable to cross-stimulation. Finally, Fig. 5(d) presents the scattering matrix for the rectangle. A striking difference here is the large number of combinations that result in stimulated generation. These results show that stimulated magnon scattering can appear in a wide variety of geometries, but its prevalence likely depends on details related to the spectrum of the normal modes and coupling efficiency to the driving fields.
Based on these results, we envisage the following avenues for future study. From the theoretical perspective, it would be interesting to explore whether the stimulated processes relate to three-magnon splitting of the off-resonantly driven uniform mode, or to four-magnon processes involving recombination of parametrically-excited modes. This could be quantified by examining coupled equations of motion for the mode amplitudes, which could be constructed from spin wave theory. [19, 20] In addition, evaluation of relevant overlap integrals would also allows us to ascertain the selection rules involved in the scattering process. [23] From the experimental perspective, the prediction of qualitatively different scattering matrices for different geometrical shapes could be tested directly using micro-focus Brillouin light scattering, which has already been employed to reveal non-degenerate parametric excitation in thin-film disks. [10] On a broader level, such stimulated processes could also be investigated for nonuniform magnetic ground states such as vortices.
@@ Acknowledgements.
We thank M. Massouras and G. de Loubens for fruitful discussions. This work was supported by the Horizon 2020 Framework Programme of the European Commission under Contract No. 899646 (k-Net).
@@
The data that support the findings of this study are openly available in Zenodo at https://doi.org/10.5281/zenodo.13711262 and at https://doi.org/10.5281/zenodo.8409685, Ref. 24 and Ref. 25.